OpenStudy (anonymous):

Which shows solutions to x + 3y = 7 if x and y must be whole numbers. (Medals:1) {(0, 5), (1, 3), (2, 1)} {(2, 1), (1, 4), (0, 7)} {(1, 2), (4, 1), (7, 0)} {(0, 5), (1, 4), (2, 1)}

OpenStudy (anonymous):

@snakeyes100

OpenStudy (anonymous):

1,2 4,1 and 7,0

OpenStudy (anonymous):

AKA the third option

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

okay when its multiple choice, the problem becomes much easier, you can substitute each value for one of the options, if the first set (0,5) doesn't work, then move on to the next option, repeat untill you find one that works, then move to the second set of values in that option, as long as the numbers keep the statement true, that they are a solution, just keep trying till you find the option that has all solutions

OpenStudy (anonymous):

mmkay thanks

OpenStudy (anonymous):

so plug in the values, 0 + 3(5) = 7 is not true, so move on to the next one, 2 + 3(1) = 7 also is not true, so move on. 1 + 3(2) = 7 IS true, so we go one coordinate set further and test 4,1 and then 7,0 since all of these keep the statement true, that set is the answer